OpenStudy (anonymous):

write the expression in simplest form. \[\sqrt{64x^3}\]

OpenStudy (anonymous):

\[\sqrt{64x^2}=\sqrt{8^2x^2x}=8x\sqrt{x}\]\]

OpenStudy (anonymous):

yes, that was a typo! should have been \[\sqrt{64x^3}\] but answer is right!

myininaya (myininaya):

\[\sqrt{64}\sqrt{x^3}=8x^\frac{3}{2}=8x^\frac{2}{2}x^\frac{1}{2}=8xx^\frac{1}{2}\]

OpenStudy (anonymous):

what is with these rational exponents?

myininaya (myininaya):

i think they are lovely

OpenStudy (anonymous):

I don't , but thanks for your help!

OpenStudy (anonymous):

yes the way you write them they are. rather unnecessary in this don't you think?

OpenStudy (anonymous):

\[\sqrt{64 x ^{3}} = \sqrt{8^{2} * x^{2}*x} = \sqrt{(8x)^2 * x} = 8x * \sqrt{x}\]